OpenStudy (anonymous):

Find the cost of papering the four walls of a room 16m long, 12m broad and 8m high, with paper 2m wide at Rs 50 per metres.

OpenStudy (mathmath333):



OpenStudy (mathmath333):

there are \(4\) walls with \(2\) walls having area =\(8\times 12\) and another \(2\) walls having \(area=16\times 8\)

OpenStudy (mathmath333):

so u have to calculate the \(\text{total area }\) of \(4\) walls

OpenStudy (mathmath333):

and then divide it by \(50\) to get the required cost.

OpenStudy (anonymous):

im getting Rs 2800

OpenStudy (mathmath333):

the total area of 4 walls is \(=2(16\times 8+8\times12)=448\)

OpenStudy (mathmath333):

r their options given ?

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

i think cost should be \(=448\times 50=22440 \) but m not sure let me call someone @AJ01

OpenStudy (anonymous):

total area is right, 8*2*(16+12)=448

OpenStudy (anonymous):

what is Rs 50?

OpenStudy (anonymous):

if this the cost per meter then you are right.....